What is the pH of 0.00028 M KOH? Is this solution acidic, basic, or neutral?
1
Expert's answer
2019-05-20T08:25:18-0400

"[KOH]=[OH-]=0.00028 mol\/L"

"[H3O+][OH-]=Kw"

"[H3O+]=Kw\/[OH-]=1* 10^{-14}\/0.00028=3.57*10^{-11}"

"pH=-log[H3O+]=-log(3.57*10^{-11})=-10.45"

It is basic solution